### v1

# Moduli spaces and breather dynamics of analytic solutions in chiral magnets without Heisenberg exchange interaction

Abstract
We investigate the special case of the chiral magnet with vanishing Heisenberg exchange energy, whose axisymmetric Skyrmion solution has previously been found. The dynamical equations of this model look like inviscid fluid flow, and by investigating path-lines of this flow we can construct explicit static and dynamic solutions. We find an infinite-dimensional family of static Skyrmions that are related to the axisymmetric Skyrmion by co-ordinate transformations thus discovering a new moduli space, and further infinite-dimensional families of axisymmetric and non-axisymmetric breather-like supercompactons. We call the maps generating such moduli space “contour-preserving” maps.
1 Introduction
The magnetic Skyrmion is a spin texture in two spatial dimensions stabilized by the chiral interaction known as Dzyaloshinskii-Moriya interaction (DMI) [1, 2] as well as an external potential in the form of the Zeeman energy or the anisotropy term [3]. The DMI is related to the spin-orbit interaction and breaks parity, energetically favouring either Skyrmions or anti-Skyrmions, but not both. The DMI term is present in chiral magnets and thin films thereof and indeed the magnetic Skyrmion has first been experimentally discovered in MnSi [4]. The DMI tends to enlarge the solitonic texture – the Skyrmion [5], whereas the potential tends to shrink it. The balance of these two forces gives the size of the Skyrmion. The Heisenberg exchange energy (relativistic kinetic energy in $\sigma$-model language) on the other hand plays no role in this balance of forces – it is classically conformally invariant in two dimensions. Nevertheless, the Heisenberg energy ensures that the soliton is completely smooth, at least in the continuum limit.
The single magnetic Skyrmion possesses radial symmetry in the sense that its energy is axisymmetric. The field is axisymmetric in the sense that as a vector field it is invariant under rotation. The phase of the Skyrmion is fixed by the chiral interaction to minimize the energy of the DMI, such that its integral is negative. Although the texture is axisymmetric and the profile is described by a simple second-order differential equation, its solution cannot be written explicitly in closed form. It is quite rare that such equations have analytic solutions, with the exception of integrable or sometimes supersymmetric models. For superconductors, the latter are known as critically coupled vortices, but even these do not possess a known analytic solution.
In the case of magnetic Skyrmions, several exact analytic solutions are known in certain limits of the general theory. The BPS (Bogomol’nyi-Prasad-Sommerfield) or supersymmetric limit possesses many solutions, including the single magnetic Skyrmion [6], but a special potential of the form of the Zeeman energy squared is required: the Zeeman energy (external magnetic field) must match exactly half of the effective anisotropy coupling, and it should also be half of the square of the DMI coupling. This precise limit is related to BPS solutions in a related gauged non-linear sigma model [7, 8]. The single magnetic Skyrmion solution was found away from this limit by Döring and Melcher, where the relation between the the coupling of this potential and DMI coupling is relaxed [9]. Finally, in a recent paper by some of the authors [10], several analytic solutions can be applied to the general theory of magnetic Skyrmions in two limits: the lump limit, where the solution is described mainly by the Heisenberg energy and the opposite limit, where the Heisenberg energy is negligible – the restricted limit. The exact solution by Döringer and Melcher is in fact of the lump type, whereas the restricted analytic solutions depend strongly on the potential in question. The restricted limit of the magnetic Skyrme theory contains only the DMI and a potential and the solutions can be compactons, supercompactons or normal solitons with a tail. Compactons are solitons with compact support, that usually have discontinuous derivatives of their fields at the compacton boundary – but a continuous energy density. Supercompactons on the other hand, are also discontinuous in the fields themselves [10].
One may think that the limit of the theory without the Heisenberg exchange interaction is rather strange. Nevertheless, it quite naturally arises in certain supersymmetric models, where such a term appears as the kinetic energy, where supersymmetrization of the theory “eats” the kinetic term [11, 12, 13]. More precisely, this happens for $\mathcal{N}=2$ supersymmetric extensions, because the solution of the auxiliary field contains the kinetic term with the opposite sign, when substituted back into the Lagrangian [^1] This also happens in $3+1$ dimensions for a Skyrme-like theory [14].. Moreover, exact solutions in a certain limit can be used to approximate solutions in the vicinity of that limit, as for example the solutions in the BPS limit were used to find novel solutions for a range of coupling parameters [15].
A property that often comes with BPS or supersymmetric solutions, is the existence of a moduli space – a space of solutions with the same energy. This is often a result of the large amount of symmetry present in such theories, severely limiting the types of viable potentials. The moduli space is not just an abstract mathematical notion, but can often be used to approximate dynamics of solitons in the near-BPS limit, by describing them as points on the moduli space instead of as solutions of full field equations. The reason why this is not trivial at all, is the fact that the moduli space is often endowed with a curved and sometimes complicated metric. The first example of this was put forward by Manton in the seminal paper [16] (see also Ref. [17]), where the scattering of BPS monopoles is described by geodesic motion on the 2-monopole moduli space.
In Ref. [10], axisymmetric analytic solutions were found for the restricted magnetic Skyrmion. Due to the limitation on the rotation of the magnetic Skyrmion coming from the DMI, it was conceivable that there would not be a nontrivial moduli space. In this paper, however, we find an infinite-dimensional moduli space. We study the equations of the restricted magnetic Skyrmion from the point of view of fluid paths and find that they are described by perfect circles. It turns out that the circles need not be concentric, since the centre can “move” as the radius changes. The only constraint is that the circles may not intersect each other. We call the symmetry connecting one such solution to another a “contour-preserving” map, and these maps generate the moduli space. We thus propose that the restricted magnetic Skyrmion has a moduli space described by the configurations of a “string” which has zero tension up to a certain maximum extension. Reintroduction of an infinitesimal Heisenberg exchange energy gives the string a tension.
This paper is organized as follows. In Sec. 2 we introduce the restricted model and study the moduli space of the analytic solutions from the point of view of fluid paths. In Sec. 3, we consider the time-dependent counterparts that resemble breather-like Skyrmions. In Sec. 4, we contemplate the extension of the model to three dimensions, but conclude that Hopfions do not exist. We conclude in Sec. 5 with a discussion and outlook. We relegate some supplemental material to three appendices.
2 Fluid flow dynamics and moduli space
Magnetic Skyrmions are described by families of topological maps $\bm{n}:\mathbb{R}^{2}\to S^{2}$, 3-dimensional magnetization vectors $\bm{n}=(n_{1},n_{2},n_{3})$, such that $|\bm{n}|^{2}=1$ with $n_{3}$ being the out-of-plane magnetization. These vortex-like configurations in chiral magnetic materials are usually described by a standard energy functional comprising a free energy Heisenberg term ($E_{2}$), the characteristic DMI term [1, 2] ($E_{1}$) and a generic external potential ($E_{0}$) such as the Zeeman interaction term. It is well-known that this model admits nontrivial topological solutions [3, 18].
In this paper, let us consider such standard two-dimensional chiral magnet energy, but in the limit where the Heisenberg exchange energy $E_{2}=\frac{1}{2}(\nabla\bm{n})^{2}$ goes to zero – the so-called restricted limit in Ref. [10], where $\nabla=(\partial_{1},\partial_{2},0)$. We hence arrive at the energy
$\begin{array}[]{rlc}E(\bm{n})&=\displaystyle\int\mathop{}\!\mathrm{d}^{2}x\Big{(}k\bm{n}\cdot(\nabla\times\bm{n})+V(\bm{n})\Big{)}\\ &\equiv E_{1}+E_{0}\ ,\end{array}$ (1)
where $k$ is the DMI strength and $V(\bm{n})$ is a generic external potential. We assume that the minimum of the energy is a point and impose the boundary condition that the field must approach this point at “large” distances: $\lim_{|\bm{x}|\to\infty}\bm{n}=\bm{N}$. This effectively point-compactifies the plane to a 2-sphere, giving the topology to the Skyrmions, $\bm{n}:S^{2}\to S^{2}$, which are characterized by the second homotopy group
$\pi_{2}(S^{2})=\mathbb{Z}\ni Q,\qquad Q=\frac{1}{4\pi}\int\mathop{}\!\mathrm{d}^{2}x\;\bm{n}\cdot\partial_{1}\bm{n}\times\partial_{2}\bm{n}\ .$ (2)
Due to the absence of the Heisenberg energy term $E_{2}$, it turns out that the assumption of a solution reaching the boundary condition is not always met. Indeed we will shortly see that there are not only compact solutions, but supercompactons, where the topological charge is fractional (see e.g. appendix. B and Ref. [10]).
The preserved rotational symmetry group of the potential energy term $E_{0}$ is the internal $\operatorname{SO}(2)_{\rm int}$, corresponding to rotations about the vacuum direction $\bm{N}$, and the spatial rotational symmetry in the $(x,y)$-plane, $\operatorname{SO}(2)_{\rm space}$. When the DMI term is included, i.e., for the total energy $E=E_{1}+E_{0}$, the independent internal and spatial rotational symmetries are broken. However, a diagonal combination of these symmetries, $\operatorname{SO}(2)_{\rm diag}$, is preserved. This symmetry corresponds to the transformation of $\bm{n}$ as a vector field under co-ordinate rotation. The separate reflections of internal and spatial reflection symmetries are broken to a single $\mathbb{Z}_{2}$ symmetry, but this one corresponds to the reflection of $(n_{1},n_{2})$ and $(x,y)$ on orthogonal axes. Taking into account these symmetries of the model, the unbroken symmetry group is at least
$G=\operatorname{O}(2)_{\rm diag}\ltimes T_{2}\ ,$ (3)
where $T_{2}$ represents two-dimensional translations, $\operatorname{O}(2)_{\rm diag}$ includes both $\operatorname{SO}(2)_{\rm diag}$ and the diagonal parity symmetry described above, and $\ltimes$ is the semi-direct product. When the energy involves the Heisenberg energy term $E_{2}$, this is the whole symmetry group. We see below that without it, the full symmetry is in fact much larger.
Let us now consider undamped Landau-Lifshitz dynamics [19] with respect to this energy:
$\partial_{t}\bm{n}=\bm{n}\times\frac{\delta E}{\delta\bm{n}}=-2k(\bm{n}\cdot\nabla)\bm{n}+\bm{n}\times\frac{\mathop{}\!\mathrm{d}V}{\mathop{}\!\mathrm{d}\bm{n}}\ .$ (4)
We can interpret this dynamical equation as a fluid flow equation with material derivative $D_{t}:=\partial_{t}+2k\bm{n}\cdot\nabla$ such that Eq. (4) reduces to
$D_{t}\bm{n}=\bm{n}\times\frac{\mathop{}\!\mathrm{d}V}{\mathop{}\!\mathrm{d}\bm{n}}\ .$ (5)
That means we can switch to a “Lagrangian” description of the Landau-Lifshitz dynamics where we imagine spins being advected by a velocity field $\vec{\nu}=(n_{1},n_{2})$ at speed $2k$, and as we follow a spin, it precesses around the effective magnetic field $\frac{\mathop{}\!\mathrm{d}V}{\mathop{}\!\mathrm{d}\bm{n}}$. Indeed we have that the “convective derivative” is equal to $2k\vec{\nu}\cdot\nabla$, being $\partial_{3}:=0$. Explicitly, we can define a path in (2+1)-dimensional spacetime, $\vec{X}(t)$, such that $\dot{\vec{X}}=2k\vec{\nu}$ and Eq. (4) reduces to
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\bm{n}(\vec{X}(t),t)=\partial_{t}\bm{n}+\dot{\vec{X}}\cdot\nabla\bm{n}=D_{t}\bm{n}=\bm{n}(\vec{X}(t),t)\times\frac{\mathop{}\!\mathrm{d}V}{\mathop{}\!\mathrm{d}\bm{n}}\bigg{\rvert}_{\bm{n}(\vec{X}(t),t)}.$ (6)
In the fluid description, $\vec{X}(t)$ is the trajectory of an infinitesimal fluid packet, also called a “path-line”. In various cases, we can solve the above equation and explicitly find this path-line and the variation of $\bm{n}$ along it. By putting together path-lines that fill every point in space and time, we can construct dynamical solutions. This is an example of the general method of characteristics [20], which is generally useful in solving first-order nonlinear equations.
For the rest of this paper, we only consider potentials that are symmetric around $\bm{e}_{3}$ (“axisymmetry”), meaning $V(\bm{n})=f(n_{3})$. In the standard cases, the potential corresponds to the combination of Zeeman interaction with strength $h$ and uniaxial anisotropy with strength $u$, i.e.
$f(n_{3})=h(1-n_{3})+u(1-n_{3}^{2})\ .$ (7)
If we define the complex velocity field $\nu:=n_{1}+\mathrm{i}n_{2}$, then we see that the evolution of a spin along the path-line $\vec{X}(t)$ is given by:
$D_{t}n_{3}=0\ ,$ (8)
$D_{t}\nu=-\mathrm{i}\nu f^{\prime}(n_{3})\ ,$ (9)
so $n_{3}(\vec{X},t)$ is constant and $\nu(\vec{X},t)=-\sqrt{1-n_{3}^{2}}e^{-\mathrm{i}f^{\prime}(n_{3})t}$. We can thus explicitly calculate the curve $\vec{X}$ introducing $X=X_{1}+\mathrm{i}X_{2}$:
$\dot{X}=-2k\sqrt{1-n_{3}^{2}}e^{-\mathrm{i}f^{\prime}(n_{3})t}\implies X(t)=X_{c}+\underbrace{\frac{-\mathrm{i}2k\sqrt{1-n_{3}^{2}}}{f^{\prime}(n_{3})}}_{\mathrm{i}R(n_{3})}e^{-\mathrm{i}f^{\prime}(n_{3})t}\ ,$ (10)
i.e.,
$X(t)=X_{c}+R(n_{3})\big{[}\sin(f^{\prime}(n_{3})t)+\mathrm{i}\cos(f^{\prime}(n_{3})t)\big{]}\ ,$ (11)
so we see that the curve is generically a circle, with radius $R$ depending on $n_{3}$. There are two special cases: if $n_{3}=\pm 1$, then the radius is zero and $\vec{X}(t)$ is a point, $X_{c}$. Notice that in the case of standard potential (7), $f(-1)>f(1)=0$ so the lowest energy state is $n_{3}=1$ everywhere. If on the other hand there are stationary points of $f(n_{3})$ for $-1<n_{3}<1$, then the radius is infinite and $\vec{X}(t)$ is a straight line.
2.1 Stationary flows
We can also look for the special cases when the flow is stationary, meaning that at each point in space, the direction of fluid flow remains the same over time. Returning to the conventional way of looking at magnetization dynamics, this means that $n_{1}$ and $n_{2}$ are static, and thus the whole magnetization configuration is static.
What we have done at this point is to construct a generic path-line for fluid flow. If we first specialize to look for stationary solutions, this path-line is a streamline: all points on this circle instantaneously have the same $n_{3}$, and $\nu$ is such that it is always tangent to the circle. By looking at $\lvert\vec{X}(t)-\vec{X}_{c}\rvert$, the circle will have a signed radius that depends on $n_{3}$, i.e.
$R(n_{3})=\frac{-2k\sqrt{1-n_{3}^{2}}}{f^{\prime}(n_{3})}\ .$ (12)
If $R>0$, it indicates a curve with spins tangent to the curve in an anticlockwise direction, conversely $R<0$ in a clockwise direction, as shown in Fig. 1.
[…]
Figure 1: “Streamlines” of constant $n_{3}$ for positive and negative $R(n_{3})$.
2.2 Static axisymmetric Skyrmion-like solution
Let us take this a step further and create an axisymmetric Skyrmion-like solution by assembling a series of concentric circles (here we set $X_{c}=0$). For this to be possible, $R(n_{3})$ must be increasing in absolute value as we start from $n_{3}=-1$ (or the circles will intersect). We will see below that in many cases $R(n_{3})$ is not monotonic all the way from $n_{3}=-1$ to $n_{3}=1$, so we must also consider discontinuous solutions where we stop before $n_{3}=1$. In this case, where our constructed solution only reaches a finite radius $R_{\text{max}}$, we limit the domain to be a disc of radius $R_{\text{max}}$, as shown in Fig. 2. Note that if $f^{\prime}(1)\neq 0$, then $R\to 0$ as $n_{3}\to 1$ and it must have reached a maximum at $n_{3}<1$ and we cannot have a whole Skyrmion.
Considering the standard potential (7) in Eq. (12), we get the dependence of radius on $n_{3}$:
$R(n_{3})=\frac{2k\sqrt{1-n_{3}^{2}}}{h+2un_{3}}\ .$ (13)
Suppose that $R$ is increasing at least until $R_{\text{Sk}}=R(0)$, the radius canonically defined as the Skyrmion radius, i.e.,
$R_{\text{Sk}}=\frac{2k}{h}\ .$ (14)
In this model, the Skyrmion radius is purely determined by the balance of DMI and Zeeman potential, and the anisotropy only determines the profile around this.
Notice that there exists a special case when $h+2u=0$, then $f^{\prime}(1)=0$ and therefore there is a chance of non-zero $R(1)$. Expanding $n_{3}=1-\frac{\epsilon^{2}}{2}$ and substituting in $f^{\prime}(n_{3})$, we find that $R\to 2k\epsilon/(u\epsilon^{2})$, so the radius is in fact infinite for $\epsilon\to 0$. Let us explicitly calculate this case: $R(n_{3})=\frac{2k\sqrt{1-n_{3}^{2}}}{h(1-n_{3})}=\frac{2k}{h}\sqrt{\frac{1+n_{3}}{1-n_{3}}}$, with $n_{3}=\cos\Theta(r)$ in spherical co-ordinates. We find that
$\Theta(r)=2\arctan\left(\frac{2k}{\sqrt{h}r}\right),$ (15)
is the Bogomol’nyi solution of Ref. [21]. This makes sense as these solutions had the property that they separately satisfied the Dirichlet and DMI+potential parts of the Euler-Lagrange equation [21, 10], rather than being the solution of a single combined Bogomol’nyi equation as at the critical coupling point [6].
Figure 2: Axisymmetric static Skyrmion-like solution by assembling a series of concentric circles for positive $R(n_{3})$.
Generally, we should first look for the poles of $R$. The condition for a pole is given by $f^{\prime}(n_{3})=0$, i.e.
$h+2un_{3}=0\ .$ (16)
This implies that in the subset of the polarized phase $h-2\lvert u\rvert>0$, there are no poles within the range $-1<n_{3}<1$, meaning we expect finite $R_{\text{max}}$ to be at $n_{3}<1$, and thus we find supercompactons with fractional topological charge. By contrast, in the canted phase $h+2u<0$ (where $\bm{N}\neq\bm{e}_{3}$) and in the remaining part of the polarized phase $h-2u<0$ we see $R\to\infty$ at $n_{3}=-\frac{h}{2u}$. We then have Skyrmions with fractional topological charge that are not supercompactons in both of these regions of the phase diagram, i.e. they fill the plane and have no discontinuity. To go further, we should look at the maximum of $R(n_{3})$:
$R^{\prime}(n_{3})=2k\frac{-(h+2un_{3})\frac{n_{3}}{\sqrt{1-n_{3}^{2}}}-\sqrt{1-n_{3}^{2}}2u}{(h+2un_{3})^{2}}=0\quad\implies\quad n_{3}^{\star}=-\frac{2u}{h}\ ,$ (17)
therefore
$R(n_{3}^{\star})=\frac{2k}{h\sqrt{1-\frac{4u^{2}}{h^{2}}}}\ .$ (18)
The second derivative of $R(n_{3})$ evaluated on $n_{3}^{\star}$ is equal to
$R^{\prime\prime}(n_{3}^{\star})=-\frac{2k}{h\left(1-\frac{4u^{2}}{h^{2}}\right)^{5/2}}\ ,$ (19)
which is negative due to $R(n_{3}^{\star})>0$, confirming that $R_{\text{max}}=R(n_{3}^{\star})$.
The energy of this configuration, according to Derrick scaling, is equal to
$E=-2\pi\int_{0}^{R_{\text{max}}}f(n_{3}(r))r\mathop{}\!\mathrm{d}r\ ,$ (20)
which is always negative, thus indicating a static configuration. Using $R_{\rm max}$ of Eq. (18) and $R$ of Eq. (13), we obtain the analytic result for the energy of the restricted Skyrmion
$E=\frac{k^{2}\pi}{4u^{2}}\left[-\frac{h^{2}-3hu+6u^{2}}{(h-2u)u^{2}}+\frac{(h^{2}-4u^{2})}{u}\left(2\arcsin\left(\frac{2u}{\sqrt{h^{2}-4u^{2}}}\right)+\log\left(1-\frac{4u^{2}}{h^{2}}\right)\right)\right],$ (21)
for $0<u<\frac{h}{2}$ and $h>0$, whereas if $u=0$, the energy simplifies to
$E=-\frac{20\pi k^{2}}{3h}\ ,$ (22)
where the maximal radius now is $R_{\rm max}=2k/h$.
This radially symmetric solution can be also reached by imposing radial symmetry and solving the associated first-order ODE, as shown in appendix B.
2.3 Infinite-dimensional family of non-axisymmetric solutions
By the exact same logic, we can find an infinite-dimensional family of solutions, by allowing the centre of the circle at each radius to be different. Reintroducing the centre $\vec{X}_{c}$, any function $\vec{X}_{c}(R)$ between $R=0$ and $R=R_{\text{max}}$ will give a valid solution (granted that $|\vec{X}_{c}^{\prime}(R)|<1$ as we will see shortly). Semi-explicitly, we can write this solution in polar-like co-ordinates adapted to this function:
$\bm{n}=\begin{pmatrix}-\sqrt{1-n_{3}(R)^{2}}\sin\varphi(\vec{X}_{c}(R))\\ \sqrt{1-n_{3}(R)^{2}}\cos\varphi(\vec{X}_{c}(R))\\ n_{3}(R)\end{pmatrix},$ (23)
where $\varphi(\vec{X}_{c})$ means azimuthal co-ordinates with $\vec{X}_{c}$ as the centre, and $R^{2}=(x_{1}-X_{c,1}(R))^{2}+(x_{2}-X_{c,2}(R))^{2}$. This equation has $R$ on both sides and is thus hard to effect as a co-ordinate transformation for a general function.
In order for circles of constant $n_{3}$ not to intersect (which will relate to the well-definedness of this co-ordinate transformation), this function must satisfy
$\lvert\vec{X}_{c}^{\prime}(R)\rvert<1\ ,\quad\forall R\in[0,R_{\text{max}}]\ .$ (24)
This gives a moduli space of static Skyrmion-like solutions described by the space
$\{\vec{X}_{c}\in C_{0}([0,R_{\text{max}}],\mathbb{R}^{2}):\lvert\vec{X}_{c}^{\prime}(R)\rvert<1\text{ where it is defined}\}\ .$ (25)
The simplest solution, $\vec{X}_{c}^{\prime}(R)=\vec{0}$, gives the axisymmetric solutions above, while in general a solution will still be contained within a circle of radius $R_{\text{max}}$ but have circles of constant $n_{3}$ at different displacements within. “Where it is defined” allows for the curve $\vec{X}_{c}(R)$ to have corners.
The boundaries of this moduli space are given by $\lvert\vec{X}_{c}^{\prime}(R)\rvert=1$ for some $R$. In the extreme case where $\vec{X}_{c}^{\prime}(R)=\vec{e}_{1}$ for all $R$, the centre of the Skyrmion touches the edge and at this point the magnetic field is not defined. In appendix A, a sanity check is presented to verify the validity of the fluid flow equations evaluated on the solutions above.
2.4 Infinite-dimensional symmetry of the energy
We check the result in a different way, that also reveals the more general symmetry of the model. We parametrize the map $\bm{n}=(\vec{\nu},n_{3})$ with $\vec{\nu}=(n_{1},n_{2})$ and $n_{3}=\sqrt{1-\nu_{i}\nu_{i}}$. Thus the energy for the restricted model can be written as
$\begin{array}[]{rlc}E(\bm{n})&=\displaystyle\int\mathop{}\!\mathrm{d}^{2}x\Big{(}2kn_{3}\epsilon_{ij}\partial_{i}\nu_{j}+V(n_{3})\Big{)}\\ &\equiv\displaystyle E_{1}+E_{0}=\int\mathop{}\!\mathrm{d}^{2}x\big{(}{\cal E}_{1}+{\cal E}_{0}\big{)}\ ,\end{array}$ (26)
where we have integrated by parts. Let us take a generic map $x_{i}\to\bm{n}$, not necessarily a solution of the equation of motion. Then consider $\gamma(n_{3})$, the preimage of $n_{3}$ constant. Let us call $a(\gamma(n_{3}))$ the area inside the curve $\gamma(n_{3})$. We then can write the potential contribution to the energy as
$E_{0}=\int\mathop{}\!\mathrm{d}n_{3}\frac{\mathop{}\!\mathrm{d}a}{\mathop{}\!\mathrm{d}n_{3}}V(n_{3})\ .$ (27)
We notice that all the maps with the same area $a(\gamma(n_{3}))$ also have the same $E_{0}$. In particular this is achieved by all maps $\Gamma:\mathbb{R}^{2}\to\mathbb{R}^{2}$ where $\gamma(n_{3})$ and $(F\circ\gamma)(n_{3})$ are connected by a rigid translation. Notice that $\Gamma$ is not necessarily an area-preserving diffeomorphism. So ${\cal E}_{0}$ is not necessarily invariant, but $E_{0}$ is invariant. Let us define the circulation of $\vec{\nu}$ around the curve $\gamma$,
$c(n_{3})=\oint_{\gamma(n_{3})}\mathop{}\!\mathrm{d}\vec{\ell}\cdot\vec{\nu}\ ,$ (28)
and by using Stokes theorem we can write the DMI term as
$E_{1}=\int kn_{3}\frac{\mathop{}\!\mathrm{d}c(n_{3})}{\mathop{}\!\mathrm{d}n_{3}}\mathop{}\!\mathrm{d}n_{3}\ .$ (29)
$E_{1}$ remains invariant if and only if $c(n_{3})$ is invariant. This is the case of rigid translations of $\gamma(n_{3})$. So the maps where $\gamma(n_{3})$ and $(\Gamma\circ\gamma)(n_{3})$ are connected by a rigid translation, are symmetries of both functionals $E_{0}$ and $E_{1}$, and consequently of $E(\bm{n})$. We may call these symmetries “contour-preserving” maps. Thus we have an $\infty$-dimensional symmetry group, $\mathcal{T}^{\rm rigid}_{2}$, of the whole energy functional (26), valid for all maps, not only the solutions of the equations of motion.
We can thus calculate the stress-energy tensor through Noether’s theorem. The total Lagrangian density $\mathcal{L}$ for this system, assuming a static configuration, is just the negative of the energy density, $\mathcal{L}=-(\mathcal{E}_{1}+V(n_{3}))$. The energy-momentum tensor associated to rigid translationally invariant symmetry, is given by
$\begin{array}[]{rlc}T_{ab}&=\displaystyle-\frac{\partial\mathcal{L}}{\partial(\partial^{a}\nu_{i})}\partial_{b}\nu_{i}+\delta_{ab}\mathcal{L}\\ \\ &=\displaystyle 2kn_{3}\epsilon_{aj}\partial_{b}\nu_{j}-\delta_{ab}(2kn_{3}\epsilon_{ck}\partial_{c}\nu_{k}+V(n_{3}))\ ,\end{array}$ (30)
where $a,b,c=1,2$ are the spatial indices. This model is not fully invariant under spatial translations due to the presence of the external potential term $V(n_{3})$. However, the model is invariant under rigid (global) translations, which means translations that shift the entire configuration of $\bm{n}$ uniformly without altering its structure relative to the co-ordinates, i.e. $\bm{n}(\vec{x})\to\bm{n}(\vec{x}-\vec{a})$.
Being $n_{3}$ fixed and constant, the system reduces effectively to a two-dimensional vector field $\nu$, subject to the constraint $\nu_{i}\nu_{i}=1-n_{3}^{2}=\mathrm{const.}$, leading to an $S^{1}$-symmetry in the $(n_{1},n_{2})$-plane. Finally, the total infinite-dimensional symmetry group of the restricted model is given by
$\mathcal{G}_{\infty}=\mathcal{T}^{\rm rigid}_{2}\ltimes\operatorname{O}(2)\ .$ (31)
2.5 Effective string-like model
As we noticed above in Eq. (25), the moduli space can be parametrized by $\vec{X}_{c}(R)$, the line of the centres depending on $R(n_{3})$ as in Eq. (13). From this interesting prospective, the moduli space is a sort of “string model” with $R$ parametrizing the string and $\vec{X}_{c}$ giving its embedding in space. This is shown pictorially in Fig. 3. In the chiral magnet model without the Heisenberg term, all the configurations (24) have the same energy; thus the effective string would have an infinite square-well elastic potential, tensionless up to some maximum extension. At first level of approximation, we can compute the effect of the Heisenberg contribution by evaluating $E_{2}$ on the solution of the model (26). We know that only the $\operatorname{SO}(2)$ invariant solutions with $\vec{X}^{\prime}_{c}(R)=\vec{0}$ will be the true minima of the energy.
The $E_{2}$ term reads
$E_{2}[\Theta,\Psi]=\int\mathop{}\!\mathrm{d}^{2}x\,\left(\frac{1}{2}(\partial_{i}\Theta)^{2}+\frac{1}{2}\sin^{2}\Theta(\partial_{i}\Psi)^{2}\right),$ (32)
with $\bm{n}=(\sin\Theta\cos\Psi,\sin\Theta\sin\Psi,\cos\Theta)$ and the fundamental state is of the form: $\Theta=\Theta(r)$ and $\Psi=\phi+\delta$, where $\delta=\frac{\pi}{2}$ for the Bloch-type DMI. Note that for the supercompacton [10], $E_{2}$ has a divergence near the boundary where $\lim_{r\to R_{\rm max}}\Theta^{\prime}(r)=\infty$. Now we perturb the solution with a modulation of the centres $\vec{X}_{c}(R)$ and evaluate the variation $\Delta E_{2}$. We use polar co-ordinates centred in each circle of radius $r$. At first order the distance between the two circles is $\delta r(\phi)=\big{(}1+|\vec{X}^{\prime}_{c}(R)|\cos(\phi-\phi_{\vec{X}_{c}})\big{)}\mathop{}\!\mathrm{d}r$ where $\phi_{\vec{X}_{c}}=\arctan({X}^{\prime}_{c,2}/{X}^{\prime}_{c,1})$. The area is in fact preserved as $\int\mathop{}\!\mathrm{d}\phi\delta r(\phi)=2\pi\mathop{}\!\mathrm{d}r$. The gradient $\partial_{i}\Psi$ is not affected by the circle’s displacement, thus $\Delta E_{2}$ gets a contribution only from the term with $\partial_{i}\Theta$. The result is then
$\Delta E_{2}=\int\mathop{}\!\mathrm{d}r\mathop{}\!\mathrm{d}\phi\;r\frac{1}{2}(\partial_{r}\Theta)^{2}\left(\frac{1}{1+|\vec{X}^{\prime}_{c}(R)|\cos(\phi-\phi_{\vec{X}_{c}})}-1\right).$ (33)
(There is one factor $\delta r$ from the measure, and one $\frac{1}{\delta r^{2}}$ from the gradient.) For small displacements,
$\Delta E_{2}=\int\mathop{}\!\mathrm{d}r\;r\frac{1}{8}(\partial_{r}\Theta)^{2}|\vec{X}^{\prime}_{c}(R)|^{2}\ .$ (34)
This is essentially a tension term for the string, where the tension density is $r\frac{1}{8}(\partial_{r}\Theta)^{2}$ which is dependent on $r$. If $R\to R_{\rm max}$ sufficiently fast, such that $\vec{X}^{\prime}_{c}(R)\to 0$, then $\Delta E_{2}$ can be finite, even for a supercompacton where $E_{2}$ diverges.
Figure 3: Pictorial illustration of the string-like model.
2.6 Explicit solutions
We will now construct a few examples of explicit solutions and numerically verify that they indeed satisfy the full equations of motion. The solutions are shown in Fig. 4.
linear quadratic trigonometric
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Three solutions of restricted magnetic Skyrmions, with different non-trivial “string motion” in the case of the Zeeman potential $h=1$, $u=0$, giving rise to a supercompacton. The four rows correspond to the collection of circles and the fields $n_{3}$, $n_{1}$ and $n_{2}$, respectively. In this figure $k=1$.
The three solutions shown in the figure correspond to the following string motions:
$\textrm{linear}:X_{c}(R)$ $=-\frac{R}{2},$ (35)
$\textrm{quadratic}:X_{c}(R)$ $=\frac{-3R^{2}}{16}+\frac{\mathrm{i}R}{4},$ (36)
$\textrm{trigonometric}:X_{c}(R)$ $=-\frac{1}{4}\cos(\pi R)+\frac{\mathrm{i}}{8}\sin(\pi R),$ (37)
where we have chosen $h=1$, $u=0$ leading to supercompactons with $R\in[0,2]$. The evaluation of the equation of motion (4) vanishes to numerical precision of the derivatives computed numerically on an interpolated field in the Mathematica package. This confirms that the collection of fluid paths indeed gives a full solution to the linear PDE.
3 Dynamics of breather-like solutions
We can also understand dynamical solutions through these characteristics, by relaxing the condition that flows are stationary. Note that from (10) we can define $\omega(n_{3})=-f^{\prime}(n_{3})$, which represents the frequency with which a single spin, followed along its trajectory, returns to the initial point.
Let us then consider the set of points where $n_{3}$ is equal to some constant. Suppose that initially this is a circle of radius $R_{0}$, with $\vec{\nu}$ tangent to the circle at all points, and that $R_{0}>R(n_{3})$. As time evolves, the curve along which $n_{3}$ is constant will contract, reaching a minimum radius $R_{\text{min}}=2R(n_{3})-R_{0}$ and then expand, reaching the radius $R_{0}$ again at time $\frac{2\pi}{\omega(n_{3})}$. If we can combine characteristics for each value of $n_{3}$, then we will produce a breathing compacton.
Systematically, we should consider a set of characteristics with equal $n_{3}$, parametrized by periodic variable $\bar{\sigma}=\frac{\sigma}{2\pi R(n_{3})}$, starting tangent to a circle of radius $R(n_{3})+\Delta(n_{3})$ at time $\Phi(n_{3})/\omega(n_{3})$ (that is, we allow different oscillations to have different phases $\Phi$). For simplicity we take $R>0$ and thus $R^{\prime}(n_{3})>0$ in this section; other possibilities are related by symmetry. Note that $\Delta$, which we can also take as positive without loss of generality, is the initial perturbation profile, it is not necessarily small. As before, we have complex co-ordinates $X$ giving the position of the characteristic and $\nu$ giving the tangent direction and speed of the characteristic with respect to time. Since we are now considering a one-parameter family of characteristics, $X$ and $\nu$ are functions of two variables, $t$ and $\bar{\sigma}$:
$X(t,\bar{\sigma})=R(n_{3})e^{\mathrm{i}\omega(n_{3})t+\mathrm{i}\bar{\sigma}}+\Delta(n_{3})e^{\mathrm{i}\bar{\sigma}+\mathrm{i}\Phi(n_{3})}\ ,\quad\quad\nu(t,\bar{\sigma})=\mathrm{i}e^{\mathrm{i}\omega(n_{3})t+\mathrm{i}\bar{\sigma}}\ .$ (38)
We see immediately that at fixed $t$, the set of points with $n_{3}$ equal is a circle centred at $X_{c}=0$ with radius
$\mathcal{R}=\sqrt{R(n_{3})^{2}+\Delta(n_{3})^{2}+2R(n_{3})\Delta(n_{3})\cos(\omega(n_{3})t-\Phi(n_{3}))}\ .$ (39)
Note that to avoid this $\mathcal{R}$-circle to not collapse to zero size, we require $\lvert\Delta\rvert<R$. To construct an extended solution, we want to put together the circles of different constant $n_{3}$ inside each other. In general, this is not possible since the period of oscillation $\frac{2\pi}{\omega(n_{3})}$ will be different for each $n_{3}$, and thus we cannot avoid the different circles colliding, at which point our field becomes discontinuous. However, with pure Zeeman interaction, that is the potential (7) with $u=0$, $\omega(n_{3})=h$ and thus all circles have the same period of oscillation. Therefore, to look for breathing solutions, we must specialize to a model with only Zeeman interaction in the potential.
Even if all circles oscillate in radius at the same time, they have different amplitudes that depend on their initial radius, so we must choose those initial radii so that the circles never collide, or shrink to zero radius. Mathematically, this means finding functions $\Delta(n_{3})$ and $\Phi(n_{3})$ such that
$\begin{cases}\Delta(-1)=0\\ \Delta<R\\ \partial_{n_{3}}\mathcal{R}^{2}>0\ ,\quad\quad 0<t<\frac{2\pi}{h}\ ,\end{cases}$ (40)
for $n_{3}$ up to some maximum which may or may not be the same as $n_{3}(R_{\text{max}})$.
The third equation simplifies to
$\Delta\Delta^{\prime}+RR^{\prime}+(R\Delta^{\prime}+R^{\prime}\Delta)\cos(ht-\Phi)-R\Delta\Phi^{\prime}\sin(ht-\Phi)>0\ .$ (41)
We should thus find the strongest constraint for all $t$, which will be
$\Delta\Delta^{\prime}+RR^{\prime}-\sqrt{(R\Delta^{\prime}+R^{\prime}\Delta)^{2}+(R\Delta\Phi^{\prime})^{2}}>0\ ,$ (42)
which we can rearrange to
$(R^{\prime 2}-\Delta^{\prime 2})(R^{2}-\Delta^{2})-R^{2}\Delta^{2}\Phi^{\prime 2}>0\ .$ (43)
We see that since $\Delta^{2}<R^{2}$, it is a necessary condition that $\Delta^{\prime 2}<R^{\prime 2}$. This then provides a constraint on the maximum size of $\Phi^{\prime}$ wherever $\Delta\neq 0$ (where it is equal to zero $\Phi$ is undefined).
This gives an infinite-dimensional moduli space of periodic axisymmetric solutions. These resemble the breathing Skyrmion solutions observed in numerics, where the radius oscillates and is coupled to rotations of the helicity, that is the angle $\delta$ in Eq. (44).
[…]
Figure 5: Constant $n_{3}$ contours at different points in time. Note that while the magnetization is always tangent to its path-line, for $ht-\Phi\neq n\pi$ it is not tangent to the circle of constant $n_{3}$, in contrast to the axisymmetric static solution.
Note that it is valid, within these constraints, to have a $\Delta$ that is supported only in a subset of $[-1,n_{3}(R_{\text{max}})]$. Physically, this corresponds to a Skyrmion that only “breathes” within a certain annulus, and it is otherwise static.
In contrast to the static case, where attainment of the inequality meant an invalid solution, here $\lvert\Delta\rvert=R$ or $\lvert\Delta^{\prime}\rvert=R^{\prime}$ at some $n_{3}$ will in general give the development of a singularity in finite time. This could occur at a point, at the centre of the Skyrmion, or along a circle.
These solutions might prove useful as starting points for understanding periodic solutions and finite-time collapse in the full model with symmetric exchange interaction.
Note that the $\operatorname{U}(1)$ action where we rotate both $\bm{n}$ and $\vec{x}$ in opposite directions is a symmetry of the energy, and this symmetry is preserved by Landau-Lifshitz dynamics. Thus we would be justified in constraining the dynamical Euler-Lagrange equations to the most general $\operatorname{U}(1)$ ansatz:
$\bm{n}=\begin{pmatrix}\sin\Theta(r)\cos(\phi+\delta(r))\\ \sin\Theta(r)\sin(\phi+\delta(r))\\ \cos\Theta(r)\end{pmatrix}\ ,$ (44)
where $(r,\phi)$ are polar co-ordinates in the plane. If $\delta(r)$ is constant, it is the helicity variable described above. Solving the resulting coupled ODEs for $\Theta(r)$ and $\delta(r)$ should then give the same set of solutions.
We can get $\delta$ from (38), more conveniently expressed as a function of $n_{3}$ (we find $n_{3}(r,t)$ by inverting $\mathcal{R}(n_{3},t)$):
$\delta(n_{3}(r,t),t)=\frac{\pi}{2}-\arg\left(1+\frac{\Delta(n_{3}(r,t))}{R(n_{3}(r,t))}e^{\mathrm{i}(\Phi(n_{3}(r,t))-ht)}\right).$ (45)
This reflects the coupling between helicity and breathing that has been observed in collective co-ordinate approximations for breathing magnetic Skyrmions [22], where an ansatz is assumed with $\delta$ independent of $r$. We can find special solutions with such $\delta$ when $\Delta/R$ and $\Phi$ are constants independent of $n_{3}$.
However yet again, we can construct non-axisymmetric breathing solutions that we would not find by this method, now allowing our families of characteristics to have a centre that continually changes with $n_{3}$:
$X(t,\bar{\sigma})=R(n_{3})e^{\mathrm{i}\omega(n_{3})t+\mathrm{i}\bar{\sigma}}+\Delta(n_{3})e^{\mathrm{i}\bar{\sigma}+\mathrm{i}\Phi(n_{3})}+X_{c}(n_{3})\ ,$ (46)
$\nu(t,\bar{\sigma})=\mathrm{i}e^{\mathrm{i}\omega(n_{3})t+\mathrm{i}\bar{\sigma}+\mathrm{i}\Phi(n_{3})}\ .$ (47)
This then gives a constraint connecting $\Delta(n_{3})$, $\Phi(n_{3})$, $X_{c}(n_{3})$ and the given function $R(n_{3})$ that unifies the axisymmetric breather (where $X_{c}^{\prime}=0$) and the non-axisymmetric static solution (where $\Delta=0$). To derive this we should modify our constraint to
$\begin{cases}\Delta(-1)=0\\ \lvert\Delta\rvert<R\\ \partial_{n_{3}}\mathcal{R}>\lvert X_{c}^{\prime}(n_{3})\rvert\ ,\quad\quad 0<t<\frac{2\pi}{h}\ .\end{cases}$ (48)
The third constraint can be rewritten as
$RR^{\prime}+\Delta\Delta^{\prime}+(R^{\prime}\Delta+R\Delta^{\prime})\cos(ht-\Phi)+R\Delta\Phi^{\prime}\sin(ht-\Phi)$ (49)
-$\mathop{-}\sqrt{R^{2}+\Delta^{2}+2R\Delta\cos(ht-\Phi)}\lvert X_{c}^{\prime}\rvert>0\ .$
However in contrast to the previous case, eliminating $t$ by finding the strongest constraint involves solving an equation of the form $\sin(x)+c_{1}\sin(2x+c_{2})=0$, which is unwieldy to write in closed form.
4 Static solutions in 3D
We can generalize to a 3D chiral magnet energy which is essentially the same
$E(\bm{n})=\int\mathop{}\!\mathrm{d}^{3}x\Big{(}k\bm{n}\cdot(\nabla\times\bm{n})+V(\bm{n})\Big{)}\ ,$ (50)
the key difference being that the DMI now includes derivatives with respect to $x_{3}$, and thus from the fluid point of view, spins are advected by the vector field $\bm{n}$ itself, rather than its projection into the plane.
The rest of the argument is very similar, but now the trajectory $\bm{X}(t)=(X_{1},X_{2},X_{3})$ as we follow a single spin is a helix, with $(X_{1},X_{2})$ as before and
$X_{3}(t)=\frac{n_{3}}{2k}t\ .$ (51)
This therefore means that a static solution in 3D must be assembled from these helices where $n_{3}$ is constant, with $n_{3}=\pm 1$ the degenerate case where the helix is a vertical line. These helices have radii given by $R(n_{3})$ as before and length scale in the $x_{3}$ direction,
$L(n_{3})=\frac{2k}{n_{3}}\ .$ (52)
In the previous case, the circles of constant $n_{3}$ had two degrees of freedom, in terms of the position of the centre. These helices have two translational degrees of freedom, two orientational degrees of freedom and one periodic degree of freedom corresponding to translation along the helix.
We can certainly arrange solutions whose 2D cross sections are identical to the static solutions above, with trivial dependence in the $x_{3}$ direction, i.e. “compacton tubes”. It may also be possible to arrange compacton tubes with some $x_{3}$ modulation.
However, it will not be possible to assemble a Hopfion solution, since it is a necessary condition that preimages of nearly all values of $n_{3}$ should form closed loops. The preimages of $n_{3}$ in any solution we assemble in the way we described above, will have no values of $n_{3}$ for which the preimages form a loop. A Hopfion may exist dynamically, however.
5 Conclusion
In this paper we consider the limit of the magnetic Skyrme theory – the theory of magnetic textures in chiral magnetic materials, where the Heisenberg exchange interaction is absent or negligible. Formulating the equations in terms of a fluid-flow equation, we find the interesting result that an enlarged moduli space emerges. The fluid lines are all circles, but due to the missing Heisenberg energy term, the circles can be moved at will – as long as they do not intersect one another. This leads to the interpretation that the moduli space is that of a “string” with the speed limit, corresponding to the circles being non-intersecting. We call the maps that generate the moduli space “contour-preserving” maps. We give a few explicit examples of exact solutions that are constructed in this way and confirm that they obey the equations of motion for the fields. We further go on and contemplate the case of time-dependent solutions that correspond to breathers. Finally, we conclude that no knotted solitons or Hopfions exist in this model.
As a future direction, one could investigate whether a moduli space exists in the case of the higher-dimensional generalization of the magnetic Skyrmion, recently put forward in Ref. [23].
It would also be interesting to study other first-order solitonic systems using the assembly of fluid lines, as used in this paper.
In the discussion of breather-like solutions, we found time-periodic solutions and solutions that develop a singularity in finite time. With the introduction of an infinitesimal Heisenberg energy term, these could be used to understand periodic solutions and singularity formation in physically realistic chiral magnets.
Acknowledgements
The work of B. B.-S. is supported by the ERC starting grant SINGinGR (Grant No. 101078061), under the European Union’s Horizon Europe program for research and innovation. The work of S. B. is supported by the INFN special research project grant “GAST” (Gauge and String Theories). S. B. G. thanks the Outstanding Talent Program of Henan University and the Ministry of Education of Henan Province for partial support. The work of S. B. G. is supported by the National Natural Science Foundation of China (Grant No. 12071111) and by the Ministry of Science and Technology of China (Grant No. G2022026021L).
Appendix A Sanity check for fluid flow equations
Let us check that these solutions are valid in a more direct way. Firstly, the static Euler-Lagrange equations in the case of the axisymmetric potential (7) reduce to
$\bm{n}\cdot\nabla n_{3}=0\ ,$
$2k\bm{n}\cdot\nabla\nu=-\mathrm{i}\nu f^{\prime}(n_{3})\ .$ (A1)
Let us consider one of these curves of constant $n_{3}$. We can give this a unit-speed parametrization respecting the orientation given by the in-plane vector, $\vec{X}(\sigma)$, i.e. $\lvert\vec{X}^{\prime}(\sigma)\rvert=1$, and $\vec{X}^{\prime}(\sigma)\parallel\bm{n}(\vec{X})$. We can then introduce the Frenet frame: $\vec{T}(\sigma)=\vec{X}^{\prime}(\sigma)$, $\vec{N}=\vec{T}^{\prime}(\sigma)/\lvert\vec{T}^{\prime}(\sigma)\rvert$ form an orthonormal pair of vectors at each point on the curve $\vec{X}$, with $\vec{T}$ tangent to the curve and $\vec{N}$ normal. The signed curvature $\kappa$ is given by the Frenet-Serret equations:
$\frac{\mathop{}\!\mathrm{d}\vec{T}}{\mathop{}\!\mathrm{d}\sigma}=\kappa\vec{N}\ ,\quad\quad\frac{\mathop{}\!\mathrm{d}\vec{N}}{\mathop{}\!\mathrm{d}\sigma}=-\kappa\vec{T}\ .$ (A2)
Note that we can only do this because $\kappa$ never vanishes for the curves in question. However, since we do have a Frenet frame, we can create a local co-ordinate system $(\sigma,\tau)$ in the vicinity of the curve, $\vec{x}=\vec{X}(\sigma)+\tau\vec{N}(\sigma)$. We can then transform the static Euler-Lagrange equations to this co-ordinate system:
$\underbrace{\bm{n}\cdot\vec{T}\partial_{\sigma}n_{3}}_{\text{vanishes}}+\bm{n}\cdot\vec{N}\partial_{\tau}n_{3}=0\ ,$ (A3)
$2k\bm{n}\cdot\vec{T}\partial_{\sigma}\nu+2k\bm{n}\cdot\vec{N}\partial_{\tau}\nu=-\mathrm{i}\nu f^{\prime}(n_{3})\ .$ (A4)
Now we substitute in our solution (23), where curves of constant $n_{3}$ are circles with $\vec{\nu}$ tangent to those curves with signed radius $R(n_{3})$ given by the formula (12). (So $\kappa$ is constant and equal to $1/R(n_{3})$.) Since $\vec{\nu}$ is tangent to the curve, the first equation is trivially satisfied, then for the second equation we apply the Frenet-Serret equations:
$\partial_{\sigma}\nu=\kappa\mathrm{i}\nu=\frac{\mathrm{i}}{R(n_{3})}\nu=\frac{-\mathrm{i}f^{\prime}(n_{3})}{2k\sqrt{1-n_{3}^{2}}}\nu=\frac{-\mathrm{i}f^{\prime}(n_{3})}{2k\bm{n}\cdot\vec{T}}\nu\ ,$ (A5)
and we see that it is satisfied also.
In general, the locally adapted co-ordinates of the Frenet frame cannot be extended to global ones. However, we have shown the static Euler-Lagrange equations are satisfied locally everywhere so this is enough for a formal validity check.
The main point is that the equation (A1) depends on the derivatives only through the combination $\bm{n}\cdot\nabla=\nu\cdot\nabla$. So, once a specific solution is found, the lines tangent to $\nu$ (circles in our case), are locally solutions irrespective to what happens in their neighbourhood. Thus the lines tangent to $\nu$ can be displaced with different translation vectors without spoiling the validity of the equation of motion.
Appendix B Radially symmetric restricted solutions
Considering the potential (7) and reparametrizing it by
$h=\tilde{h}\sin\alpha\ ,\qquad u=\frac{\tilde{h}}{2}\cos\alpha\ .$ (A6)
we have that
$-f^{\prime}(n_{3})=\tilde{h}(\sin\alpha+n_{3}\cos\alpha)\ .$ (A7)
It is important that the potential only depends on $n_{3}$, otherwise further symmetry breaking will occur. For $40\leq\alpha<\pi/$ 4 we have a metastable point at $n_{3}=-1$. For $3\pi/4<\alpha<\pi$ we have spontaneous symmetry breaking, and $\alpha=3\pi/4$ is the line of the $(1-n_{3})^{2}$ potential that was considered in Ref. [21]. $\alpha=\pi/2$ corresponds to the Zeeman potential. We can therefore limit ourselves, without loss of generality, to $0\leq\alpha\leq\pi$, and $\tilde{h}\geq 0$.
We consider now the simplified case where $X_{c}:=0$, so that the equation of motion is simply given by Eq. (12) with $R=r$ being the radius in the 2-plane $r=\sqrt{x^{2}+y^{2}}$. Using $-f^{\prime}(n_{3})$ of Eq. (A7), we obtain
$r=\frac{2\ell\sin\Theta}{\sin\alpha+\cos\alpha\cos\Theta}\ ,$ (A8)
[…]
$\alpha=\pi/2\,:\quad\Theta(r)=\pi-\arcsin\rho(r)\ ,$ $\rho(r)$ $=\frac{r}{2\ell}\ ,$ (A9)
$\alpha=3\pi/4\,:\quad\Theta(r)=2\arctan\frac{1}{\rho(r)}\ ,$ $\rho(r)$ $=\frac{r}{2\sqrt{2}\ell}\ ,$ (A10)
with $\rho=\rho(r)\in[0,1)$. In general what happens is the following. For $40\leq\alpha\leq\pi/$ 4, the solutions exist on the entire range $\rho\in[0,\infty]$, but none of them reach $\Theta(\infty)\to 0$ and as a result they are fraction solutions with topological charges $Q\leq\frac{1}{2}$.
For $\pi/4<\alpha<3\pi/4$ the situation is similar to $\alpha=\pi/2$, we have a supercompacton as restricted solution, but again the topological charge is $10<Q<$ 1. As $\alpha\to 3\pi/4$ the supercompacton becomes a lump and the topological charge becomes integer, $Q=1$. The generic solution for $0\leq\alpha\leq\pi$ is
$\Theta(r)=\arccos\left(-\frac{2\sqrt{4+\rho^{2}\cos 2\alpha}+\frac{1}{2}\rho^{2}\sin 2\alpha}{4+\rho^{2}\cos^{2}\alpha}\right),\qquad\rho=\frac{r}{\ell}\ ,$ (A11)
with
$\rho\in\begin{cases}[0,\infty]\ ,&0\leq\alpha\leq\frac{\pi}{4}\ ,\\ \left[0,2\sqrt{\frac{-1}{\cos 2\alpha}}\right]\ ,&\frac{\pi}{4}<\alpha<\frac{3\pi}{4}\ ,\\ [0,\infty]\ ,&\frac{3\pi}{4}\leq\alpha\leq\pi\ .\end{cases}$ (A12)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Restricted solutions of magnetic Skyrmions with the potential (7), parametrized by Eq. (A6) with $\alpha$ in the range $0\leq\alpha\leq\pi$.
In the range $3\pi/4<\alpha\leq\pi$ we have again a fracton. The fractons that exist for $40\leq\alpha\leq\pi/$ 4 and for $3\pi/4<\alpha\leq\pi$ exist on the entire range $\rho\in[0,\infty]$ and have the asymptotic behaviour
$\lim_{r\to\infty}\Theta(r)=\arccos(-\tan\alpha)\ ,$ (A13)
which is the true vacuum of the potential. Notice that as soon as we introduce the $E_{2}$ term, this solution acquires a “log” divergent energy due to the angular derivative. So it is like a global vortex solution. The fracton has a definite topological charge and Eq. (2) reduces to
$-\frac{1}{2}\int_{\pi}^{\Theta(\infty)}\sin(\Theta)\mathop{}\!\mathrm{d}\Theta=-\frac{1}{2}[\cos\Theta]_{\Theta(\infty)}^{\pi}=\frac{1}{2}(1+\tan\alpha)\ ,\qquad\alpha\in\big{\{}[0,\pi/4],(3\pi/4,\pi]\big{\}}\ ,$ (A14)
which becomes $1/2$ for $\alpha=0,\pi$; in this case we can call it a meron.
Appendix C The restricted model in complex co-ordinates
Let us consider the static energy density (1). We now introduce complex disk co-ordinates, such that $\nu=n_{1}+\mathrm{i}n_{2}$, $\nu\bar{\nu}\leq 1$ and $n_{3}=\pm\sqrt{1-\nu\bar{\nu}}$, together with $\pm$ covering the north and south part of the 2-sphere. The derivative of the standard potential (7) is
$-f^{\prime}(n_{3})=h+2un_{3}\ .$ (A15)
Inserting the complex co-ordinates and changing the co-ordinates of the plane to $z=x^{1}+\mathrm{i}x^{2}$ with its complex conjugate, we arrive at $E=\int\mathop{}\!\mathrm{d}^{2}x\;\mathcal{E}$:
=±$\mathcal{E}=\pm\mathrm{i}k\frac{2\bar{\partial}\bar{\nu}-2\partial\nu+\bar{\nu}^{2}\bar{\partial}\nu-\nu^{2}\partial\bar{\nu}+\nu\bar{\nu}(-\bar{\partial}\bar{\nu}+\partial\nu)}{2\sqrt{1-\nu\bar{\nu}}}+f(\pm\sqrt{1-\nu\bar{\nu}})\ ,$ (A16)
with the lower sign corresponding to $n_{3}\leq 0$ and the upper sign to $n_{3}\geq 0$. For the supercompacton, the lower sign is sufficient as the other patch of the 2-sphere does not exist. The energy is real, so it suffices to write the equation of motion for $\bar{\nu}$:
$\left(2\mathrm{i}\nu\partial+2\mathrm{i}\bar{\nu}\bar{\partial}-\frac{f^{\prime}}{k}\right)\nu=0\ ,$ (A17)
which is indeed very elegant and independent of the sign $\pm$ choosing the patch. A simple solution that covers the north patch, suitable for the Zeeman potential which has $f^{\prime}=-h$, is given by
$\nu=\nu(z)=\frac{h}{2k\mathrm{i}}z\ .$ (A18)
Including $u\neq 0$ introduces a square root into the equation, making it more cumbersome.